OpenStudy (anonymous):

remainder theorum Consider the polynomial P(x)=(kx^3)+(2x^2)-5x+4. Find the value of k such that the remainder is -14 when P(x) is divided by x-2 k=?

OpenStudy (anonymous):

P(x)=(x-2)Q(x)-14

OpenStudy (anonymous):

not sure how to set this one up

OpenStudy (anonymous):

Have you done the remainder theorem before?

OpenStudy (anonymous):

somewhat

OpenStudy (anonymous):

its like 13/5=2 r 3 then 13=5*2+3

OpenStudy (anonymous):

P(x)/(x-a)=q(x)+r

OpenStudy (anonymous):

then P(x)=(x-a)q(x)+r

OpenStudy (anonymous):

Yeah, also--this is key--when you plug in the "a" to get "P(a)", the answer is "r"

OpenStudy (anonymous):

So calculate "P(2)" and tell me what you get

OpenStudy (anonymous):

its not 2

OpenStudy (anonymous):

its -2

OpenStudy (anonymous):

No, no, what I mean is this: Take your initial polynomial and plug in "2" for x and tell me the resulting expression.

OpenStudy (anonymous):

i got it thanks

OpenStudy (anonymous):

But you mentioned "x-a", right? Well, if you have "x-2" then that means that "a" is equal to "2"

OpenStudy (anonymous):

(x-(-2))

OpenStudy (anonymous):

k(2)^3+2(2)^2-5(2)+4=-14

OpenStudy (anonymous):

But in your initial question you said you were dividing by "(x-2)"?

OpenStudy (anonymous):

ok, now simplify...

OpenStudy (anonymous):

8k+8-10+4=14 8k+2=-14 8k=-16 k=-2

OpenStudy (anonymous):

i just need to keep practicing it

OpenStudy (anonymous):

Ok, ok. You're right -- k=-2. When you said "it's -2" i didn't know what you were referring to.

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

so the polynomial is equal to: -2k^3 + 2x^2 ... and do on

OpenStudy (anonymous):

thanks for your time @Splash_Dance

OpenStudy (anonymous):

Yeah, that was good--you got the answer fast!

OpenStudy (anonymous):

Good luck with the rest of your problems